# QMambaBSR: Burst Image Super-Resolution with Query State Space Model

### v1

@@ Abstract
Burst super-resolution (BurstSR) aims to reconstruct high-resolution images with higher quality and richer details by fusing the sub-pixel information from multiple burst low-resolution frames. In BusrtSR, the key challenge lies in extracting the base frame’s content complementary sub-pixel details while simultaneously suppressing high-frequency noise disturbance. Existing methods attempt to extract sub-pixels by modeling inter-frame relationships frame by frame while overlooking the mutual correlations among multi-current frames and neglecting the intra-frame interactions, leading to inaccurate and noisy sub-pixels for base frame super-resolution. Further, existing methods mainly employ static upsampling with fixed parameters to improve spatial resolution for all scenes, failing to perceive the sub-pixel distribution difference across multiple frames and cannot balance the fusion weights of different frames, resulting in over-smoothed details and artifacts. To address these limitations, we introduce a novel Query Mamba Burst Super-Resolution (QMambaBSR) network, which incorporates a Query State Space Model (QSSM) and Adaptive Up-sampling module (AdaUp). Specifically, based on the observation that sub-pixels have consistent spatial distribution while random noise is inconsistently distributed, a novel QSSM is proposed to efficiently extract sub-pixels through inter-frame querying and intra-frame scanning, while mitigating noise interference in a single step. Moreover, AdaUp is designed to dynamically adjust the upsampling kernel based on the spatial distribution of multi-frame sub-pixel information in the different burst scenes, thereby facilitating the reconstruction of the spatial arrangement of high-resolution details. Extensive experiments on four popular synthetic and real-world benchmarks demonstrate that our method achieves a new state-of-the-art performance. The code will be publicly available.
Introduction
In recent years, with the continuous development of smartphones, overcoming the limitations of smartphone sensors and lenses to reconstruct high-quality, high-resolution (HR) images has become a research hotspot. Benefited from the development of deep learning, single image super-resolution (SISR) (Ju, Schaefer, and Warren 2023; Dong et al. 2015; Yang et al. 2010; Zhang et al. 2018; Peng et al. 2024a, b) has achieved remarkable progress, but the performance is still limited by the finite information provided by a single image. Consequently, numerous researchers are dedicating their efforts to burst super-resolution (BurstSR), which aims to leverage the rich sub-pixel details provided by a sequence of burst RAW/RGB low-resolution images captured by hand-tremor and camera/object motions to overcome the limitations of SISR, achieving substantial advancements (Li et al. 2018; Chen et al. 2023; Saharia et al. 2022; Wang et al. 2018; Liang et al. 2021).
@@ Introduction
In BurstSR, the first RAW/RGB image is the super-resolution frame, denoted as the base frame, while the remaining images, referred to as current frames, supply sub-pixel information for producing a high-quality HR image. The pipeline of most existing BurstSR approaches can be mainly categorized: alignment, fusion, and upsampling. Firstly, due to the misalignment caused by hand-tremor, alignment methods (Bhat et al. 2021a; Wei et al. 2023; Dudhane et al. 2022; Luo et al. 2022) are employed to align the current frames with the target base frame. Then, the primary challenge lies in extracting sub-pixel information from the current frames that match the content of the base frame while concurrently suppressing high-frequency random noise. Previous methods, such as weighted-based multi-frame fusion (Bhat et al. 2021a, b), obtain residuals by subtracting each current frame from the base frame and utilizing simple weighting techniques to fuse obtained residual information. Although easy to perform, these methods neglect the inter-frame relationship among multi-frames, failing to extract sub-pixels that better match the base frame and are susceptible to interference from noise in RAW images. To enhance inter-frame relationships, BIPNet (Dudhane et al. 2022) proposes channel shuffling of multi-frame features to improve information flow between different frames. Consequently, recent state-of-the-art methods, such as GMTNet  (Dudhane et al. 2023; Mehta et al. 2023; Luo et al. 2021), propose using cross-attention, explicitly employing the base frame as a query to retrieve and capture feature differences from the current frame pixel-to-pixel to extract sub-pixel information. RBSR  (Wu et al. 2023) utilize RNNs (Liang and Hu 2015) for frame-by-frame feature fusion, as shown in Fig. 1 (a). However, the aforementioned methods, which extract sub-pixel information in a frame-by-frame manner, are unable to accurately and effectively extract sub-pixels to supplement the base frame. Additionally, the severe noise contained in RAW images further complicates the ability of existing methods to distinguish between sub-pixels and noise, resulting in introducing artifacts and over-smoothing details.
@@
After fusion, adaptively learning high-resolution mappings from the extracted and fused features remains a paramount challenge in BurstSR. The existing state-of-the-art methods, such as Burstormer  (Dudhane et al. 2023) and BIPNet (Dudhane et al. 2022), primarily utilize static interpolation, transposed convolution (Gao et al. 2019), or pixel shuffle for static upsampling. Nevertheless, these approaches make it difficult to adaptively perceive the variations in sub-pixel distribution across different scenes by employing static upsampling methods, resulting in the inability to utilize the spatial arrangement of sub-pixels to accurately reconstruct high-quality, high-resolution (HR) images.
@@
To address these issues, we propose a novel Query Mamba Burst Super-Resolution (QMambaBSR) network, which integrates a novel Query State Space Model (QSSM) and an Adaptive Up-sampling module (AdaUp) to reconstruct high-quality high-resolution images from burst low-resolution images. Specifically, QSSM is first proposed to efficiently extract the sub-pixels in both inter-frame and intra-frame while mitigating noise interference. In particular, QSSM retrieves information across current frames for the base frame by modifying control matrix $B$ and discretization step size $\Delta$ in the state space function, as shown in Fig. 1 (b). AdaUp is proposed to perceive the spatial distribution of sub-pixel information and subsequently adaptively adjust the upsampling kernel to enhance the reconstruction of high-quality HR images across diverse burst LR scenarios. Furthermore, to comprehensively fuse sub-pixels with different scales, the Multi-scale Fusion Module is proposed to combine channel Transformer and local CNN, as well as horizontal and vertical global Mamba, to fuse sub-pixel information of different scales. Extensive experiments on four popular synthetic and real-world benchmarks demonstrate that our method achieves a new state-of-the-art, delivering superior visual results.
@@
  • We propose a novel Adaptive Up-sampling module and a Multi-scale Fusion Module, designed respectively for adaptive up-sampling based on the spatial arrangement of sub-pixel information in various burst LR scenarios, and for the fusion of sub-pixels across different scales.
@@
In this section, we briefly review Multi-Frame Super-Resolution and State Space Models. More comprehensive surveys are provided in  (Xu et al. 2024; Bhat et al. 2022).
@@
Multi-Frame Super-Resolution. With the rapid development of deep learning in recent years (He et al. 2016; Liu et al. 2021; Vaswani et al. 2017; Goodfellow et al. 2020), deep-learning-based single image super-resolution (SISR) achieves significant breakthroughs (Yang et al. 2020; Wu et al. 2024; Chen et al. 2024b; Yue, Wang, and Loy 2024). However, due to the limited information provided by a single image, the performance of SISR is significantly constrained (Lu et al. 2022; Zhang, Gool, and Timofte 2020; Wang et al. 2024a; Simonyan and Zisserman 2014). Therefore, Multi-Frame Super-Resolution (MFSR) is proposed to overcome the limitations of SISR by leveraging the useful sub-pixel information contained in multiple low-resolution images, achieving superior high-resolution reconstruction. In particular, DBSR (Bhat et al. 2021a) proposes using optical flow methods to explicitly align multiple low-resolution images and then fuse their features through attention weights. MFIR (Bhat et al. 2021b) utilizes optical flow for feature warping and proposes a deep reparametrization of the classical MAP formulation for multi-frame image restoration. BIPNet (Dudhane et al. 2022) proposes a pseudo-burst fusion strategy by fusing temporal features channel-by-channel, enabling frequent inter-frame interaction. Burstormer (Dudhane et al. 2023) leverages multi-scale local and non-local features for alignment and employs neighborhood interaction for further inter-frame feature fusion. RBSR (Wu et al. 2023) utilizes recurrent neural networks for progressive feature aggregation. However, most of these methods mainly use frame-by-frame approaches or pairwise interactions, either failing to explicitly extract sub-pixel information from the current frames or only querying the current frame point-by-point from the base frame. This makes it difficult for them to effectively extract sub-pixel details while suppressing noise interference. To address these limitations, we propose Query Mamba Burst Super-Resolution (QMambaBSR), which allows the base frame to simultaneously query inter-frame and intra-frame information to extract sub-pixel details embedded in structured regions while also suppressing noise interference.
@@
State Space Models. State Space Models (SSMs) originated in the 1960s in control systems (Kalman 1960), where they are used for modeling continuous signal input systems. Recently, advancements in SSMs have led to their application in computer vision (Zhu et al. 2024; Patro and Agneeswaran 2024; Fu et al. 2024; Chen et al. 2024a). Notably, Visual Mamba introduced a residual VSS module and developed four scanning directions for visual images, achieving superior performance compared to ViT (Dosovitskiy et al. 2020) while maintaining lower model complexity, thereby attracting significant attention (Guo et al. 2024; Qiao et al. 2024; Tang et al. 2024; Wang et al. 2024b; Zhen, Hu, and Feng 2024; Li et al. 2024). QueryMamba (Zhong et al. 2024) is proposed to apply SSM to video action forecasting tasks. MambaIR (Guo et al. 2024) is the first to employ SSMs in image restoration, enhancing efficiency and global perceptual capabilities. However, there remains potential for further exploration of SSMs in Burst SR. Therefore, we propose a novel Query-based State Space Model designed to efficiently extract sub-pixel information for Burst SR.
Method
Overview
@@
Given a sequence of input RAW/RGB low-resolution (LR) images, denoted as $\{x_{i}\}_{i=1}^{N}$, where $N$ represents the number of burst LR frames. Following (Bhat et al. 2021a), we denote the first image as the base frame for super-resolution, while the other LR frames are used to provide rich sub-pixel information and are referred to as current frames. BurstSR can be defined as utilizing the sub-pixel information extracted from the current frames to supplement the base frame, generating a high-quality, high-resolution RGB image $I_{HR}$ with a super-resolution factor of $s$.
@@
To achieve this goal, we propose QMambaBSR for burst image super-resolution, as illustrated in Fig. 2. First, we use alignment block (Dudhane et al. 2023) to align the current images to the spatial position of the base frame. Next, we introduce a novel Query State Space Model (QSSM) designed to query sub-pixel information from the current images and mitigate noise interference in both inter-frame and intra-frame manner. Additionally, we present a novel Adaptive Up-sampling (AdaUp) module, which facilitates adaptive up-sampling based on the spatial arrangement of sub-pixel information in various burst images. Finally, a new Multi-scale Fusion Module is incorporated to fuse sub-pixel information across different scales. Next, we provide a detailed explanation of each component.
Query State Space Model
@@
Considering that burst RAW images often contain high-frequency random noise and the sub-pixel information to be extracted typically shares a similar distribution with the base frame, it is crucial for the BurstSR task to utilize the base frame to uncover the rich sub-pixel information contained in the current frames for super-resolution, while simultaneously suppressing noise. Existing methods (Bhat et al. 2021a; Dudhane et al. 2023; Wei et al. 2023; Wu et al. 2023; Dudhane et al. 2022) simply concatenate multi-frame information but fail to precisely extract sub-pixel information from the current frames using the base frame, resulting in a scarcity of sub-pixel details and consequently making it difficult to reconstruct fine details. Moreover, while some existing approaches attempt to use cross-attention (Mehta et al. 2023), utilizing the base frame as the query for sub-pixels extraction, such frame-by-frame methods struggle to suppress noise interference and are plagued by high computational complexity. This often results in the presence of noise and the introduction of artifacts. Therefore, we propose the Query State Space Model (QSSM), which enables the base frame to efficiently query all current frames simultaneously in both intra-frame and inter-frame manners. By leveraging the consistent distribution of sub-pixels and the inconsistent distribution of noise, our QSSM can simultaneously query multiple images to extract the necessary sub-pixel information while effectively suppressing random noise.
First, let’s briefly review the State Space Model (SSM). The latest advances in structured state space sequence models (S4) are largely inspired by continuous linear time-invariant (LTI) systems, which map input $x(t)$ to output $y(t)$ through an implicit latent state $h(t)\in\mathbb{R}^{N}$  (Guo et al. 2024). This system can be represented as a linear ordinary differential equation (ODE):
@@
At this point, the LTI system’s input parameter matrices are static. Therefore, recent work (Guo et al. 2024) makes $B$, $C$, and $\Delta$ depend on the input. Recent research suggests that since the $A$ matrix is a HIPPO matrix and $\Delta$ represents the step size, $\exp(\Delta A)$ can be viewed as forget gate and input gate (Han et al. 2024), which modulates the influence of the input on the state.
However, traditional SSM lacks the multi-frame querying capabilities that are crucial for BurstSR tasks. Therefore, we propose a QSSM to enable the base frame to gate the output of the current frames, thereby allowing the base frame to perform information queries on the current frames to obtain the sub-pixels while eliminating noise, as illustrated in Fig. 1. Specifically, we let the current frames drive the state changes, with the base frame gating the influence of the current frames on the state through $B$ and $\Delta$. As shown in Fig. 2 (a), the corresponding formulas are as follows:
@@
The base frame is transformed through a learnable linear layer to generate $\Delta_{\text{base}_{t}}$ and $B_{\text{base}_{t}}$, and these are then used in the discretization formula from Eq. (2) to obtain $\overline{A}_{\text{base}_{t}}$ and $\overline{B}_{\text{base}_{t}}$. The current frames are processed through another linear layer to obtain $x_{\text{cur}_{t}}$, where $t$ indicates the positional relationship after flattening the base frame and current frames.

### v2

@@ Abstract
Burst super-resolution (BurstSR) aims to reconstruct high-resolution images by fusing subpixel details from multiple low-resolution burst frames. The primary challenge lies in effectively extracting useful information while mitigating the impact of high-frequency noise. Most existing methods rely on frame-by-frame fusion, which often struggles to distinguish informative subpixels from noise, leading to suboptimal performance. To address these limitations, we introduce a novel Query Mamba Burst Super-Resolution (QMambaBSR) network. Specifically, we observe that sub-pixels have consistent spatial distribution while noise appears randomly. Considering the entire burst sequence during fusion allows for more reliable extraction of consistent subpixels and better suppression of noise outliers. Based on this, a Query State Space Model (QSSM) is proposed for both inter-frame querying and intra-frame scanning, enabling a more efficient fusion of useful subpixels. Additionally, to overcome the limitations of static upsampling methods that often result in over-smoothing, we propose an Adaptive Upsampling (AdaUp) module that dynamically adjusts the upsampling kernel to suit the characteristics of different burst scenes, achieving superior detail reconstruction. Extensive experiments on four benchmark datasets—spanning both synthetic and real-world images—demonstrate that QMambaBSR outperforms existing state-of-the-art methods.
1 Introduction
In recent years, with the continuous development of smartphones, overcoming the limitations of smartphone sensors and lenses to reconstruct high-quality, high-resolution (HR) images has become a research hotspot. Benefited from the development of deep learning, single image super-resolution (SISR) [19, 9, 50, 53, 33, 34] has achieved remarkable progress, but the performance is still limited by the finite information provided by a single image. Consequently, numerous researchers are dedicating their efforts to burst super-resolution (BurstSR), which aims to leverage the rich sub-pixel details provided by a sequence of burst low-resolution images captured by hand-tremor and camera/object motions to overcome the limitations of SISR, achieving substantial advancements [23, 7, 36, 43, 24].
@@ 1 Introduction
In BurstSR, the first image is the super-resolution frame, denoted as the base frame, while the remaining images, referred to as current frames, supply sub-pixel information for producing a high-quality HR image. The pipeline of most existing BurstSR approaches can be mainly categorized: alignment, fusion, and upsampling. Firstly, due to the misalignment caused by hand-tremor, alignment methods [1, 45, 11, 29] are employed to align the current frames with the target base frame. Then, the primary challenge lies in extracting sub-pixel information from the current frames that match the content of the base frame while concurrently suppressing high-frequency random noise. Previous methods, such as weighted-based multi-frame fusion [1, 2], obtain residuals by subtracting each current frame from the base frame and utilizing simple weighting techniques to fuse obtained residual information. Although easy to perform, these methods neglect the inter-frame relationship among multi-frames, failing to extract sub-pixels that better match the base frame and are susceptible to interference from noise in RAW images. To enhance inter-frame relationships, BIPNet [11] proposes channel shuffling of multi-frame features to improve information flow between different frames. Consequently, recent state-of-the-art methods, such as GMTNet  [12, 31, 28], propose using cross-attention, explicitly employing the base frame as a query to retrieve and capture feature differences from the current frame pixel-to-pixel to extract sub-pixel information. RBSR  [46] utilize RNNs [25] for frame-by-frame feature fusion, as shown in Figure 1 (a). The aforementioned methods extract sub-pixel information and denoise in a frame-by-frame manner or implement sub-pixel extraction and denoising through separate modules. However, these methods overlook the distinct characteristics between sub-pixels and noise, making it challenging to capture useful sub-pixels from noise and leading to artifacts and over-smoothed details.
@@
After fusion, adaptively learning high-resolution mappings from the extracted and fused features remains a paramount challenge in BurstSR. The existing state-of-the-art methods, such as Burstormer  [12] and BIPNet [11], primarily utilize static interpolation, transposed convolution [14], or pixel shuffle [38] for static upsampling. Nevertheless, these approaches make it difficult to adaptively perceive the variations in sub-pixel distribution across different scenes by employing static upsampling methods, resulting in the inability to utilize the spatial arrangement of sub-pixels to accurately reconstruct high-quality, high-resolution (HR) images.
@@
To address these issues, we propose a novel Query Mamba Burst Super-Resolution (QMambaBSR) network, which integrates a novel Query State Space Model (QSSM) and an Adaptive Up-sampling module (AdaUp) to reconstruct high-quality high-resolution images from burst low-resolution images. Specifically, QSSM is first proposed to efficiently extract the sub-pixels in both inter-frame and intra-frame while mitigating noise interference. In particular, QSSM retrieves information across current frames for the base frame by modifying control matrix $B$ and discretization step size $\Delta$ in the state space function, as shown in Figure 1 (b). QSSM synchronously performs inter-frame querying and intra-frame scanning, comprehensively considering both inter-frame and intra-frame information to achieve integrated sub-pixel extraction and noise suppression. AdaUp is proposed to perceive the spatial distribution of sub-pixel information and subsequently adaptively adjust the upsampling kernel to enhance the reconstruction of high-quality HR images across diverse burst LR scenarios. Furthermore, to comprehensively fuse sub-pixels with different scales, the Multi-scale Fusion Module is proposed to combine channel Transformer and local CNN, as well as horizontal and vertical global Mamba, to fuse sub-pixel information of different scales. Extensive experiments on four popular synthetic and real-world benchmarks demonstrate that our method achieves a new state-of-the-art, delivering superior visual results.
@@
  • We propose a novel Adaptive Up-sampling module, designed respectively for adaptive up-sampling based on the spatial arrangement of sub-pixel information in various burst LR scenarios, and for the fusion of sub-pixels across different scales.
@@ 2 Related Work
In this section, we briefly review Multi-Frame Super-Resolution and State Space Models. More comprehensive surveys are provided in  [48, 3].
@@
Multi-Frame Super-Resolution. With the rapid development of deep learning in recent years [18, 26, 41, 15], deep-learning-based single image super-resolution (SISR) achieves significant breakthroughs [49, 47, 6, 51]. However, due to the limited information provided by a single image, the performance of SISR is significantly constrained [27, 52, 42, 39]. Therefore, Multi-Frame Super-Resolution (MFSR) is proposed to overcome the limitations of SISR by leveraging the useful sub-pixel information contained in multiple low-resolution images, achieving superior high-resolution reconstruction. In particular, DBSR [1] proposes using optical flow methods to explicitly align multiple low-resolution images and then fuse their features through attention weights. MFIR [2] utilizes optical flow for feature warping and proposes a deep reparametrization of the classical MAP formulation for multi-frame image restoration. BIPNet [11] proposes a pseudo-burst fusion strategy by fusing temporal features channel-by-channel, enabling frequent inter-frame interaction. Burstormer [12] leverages multi-scale local and non-local features for alignment and employs neighborhood interaction for further inter-frame feature fusion. RBSR [46] utilizes recurrent neural networks for progressive feature aggregation. However, most of these methods mainly use frame-by-frame approaches or pairwise interactions, either failing to explicitly extract sub-pixel information from the current frames or only querying the current frame point-by-point from the base frame. This makes it difficult for them to effectively extract sub-pixel details while suppressing noise interference. To address these limitations, we propose Query Mamba Burst Super-Resolution (QMambaBSR), which allows the base frame to simultaneously query inter-frame and intra-frame information to extract sub-pixel details embedded in structured regions while also suppressing noise interference.
@@
State Space Models. State Space Models (SSMs) originated in the 1960s in control systems [20], where they are used for modeling continuous signal input systems. Recently, advancements in SSMs have led to their application in computer vision [56, 32, 13, 5]. Notably, Visual Mamba introduced a residual VSS module and developed four scanning directions for visual images, achieving superior performance compared to ViT [10] while maintaining lower model complexity, thereby attracting significant attention [16, 35, 40, 44, 54, 22]. QueryMamba [55] is proposed to apply SSM to video action forecasting tasks. MambaIR [16] is the first to employ SSMs in image restoration, enhancing efficiency and global perceptual capabilities. However, there remains potential for further exploration of SSMs in BurstSR. Therefore, we propose a novel Query-based State Space Model designed to efficiently extract sub-pixel information for BurstSR.
3 Method
3.1 Overview
@@
Given a sequence of input low-resolution (LR) images, denoted as $\{x_{i}\}_{i=1}^{N}$, where $N$ represents the number of burst LR frames. Following [1], we denote the first image as the base frame for super-resolution, while the other LR frames are used to provide rich sub-pixel information and are referred to as current frames. BurstSR can be defined as utilizing the sub-pixel information extracted from the current frames to supplement the base frame, generating a high-quality, high-resolution RGB image $I_{HR}$ with a super-resolution factor of $s$.
@@
To achieve this goal, we propose QMambaBSR for burst image super-resolution, as illustrated in Figure 2. First, we use alignment block [12] to align the current images to the spatial position of the base frame. Next, we introduce a novel Query State Space Model (QSSM) designed to query sub-pixel information from the current images and mitigate noise interference in both inter-frame and intra-frame manner. Additionally, we present a novel Adaptive Up-sampling (AdaUp) module, which facilitates adaptive up-sampling based on the spatial arrangement of sub-pixel information in various burst images. Finally, a new Multi-scale Fusion Module is incorporated to fuse sub-pixel information across different scales. Next, we provide a detailed explanation of each component.
3.2 Query State Space Model
@@
Considering that burst RAW images often contain high-frequency random noise and the sub-pixel information to be extracted typically shares a similar distribution with the base frame, it is crucial for the BurstSR task to utilize the base frame to uncover the rich sub-pixel information contained in the current frames for super-resolution, while simultaneously suppressing noise. Existing methods [1, 12, 45, 46, 11] simply concatenate multi-frame information but fail to precisely extract sub-pixel information from the current frames using the base frame, resulting in a scarcity of sub-pixel details and consequently making it difficult to reconstruct fine details. Moreover, while some existing approaches attempt to use cross-attention [31], utilizing the base frame as the query for sub-pixels extraction, such frame-by-frame methods struggle to suppress noise interference and are plagued by high computational complexity. This often results in the presence of noise and the introduction of artifacts. Therefore, we propose the Query State Space Model (QSSM), which enables the base frame to efficiently query all current frames simultaneously in both intra-frame and inter-frame manners. By leveraging the consistent distribution of sub-pixels and the inconsistent distribution of noise, our QSSM can simultaneously query multiple images to extract the necessary sub-pixel information while effectively suppressing random noise.
First, let’s briefly review the State Space Model (SSM). The latest advances in structured state space sequence models (S4) are largely inspired by continuous linear time-invariant (LTI) systems, which map input $x(t)$ to output $y(t)$ through an implicit latent state $h(t)\in\mathbb{R}^{N}$  [16]. This system can be represented as a linear ordinary differential equation (ODE):
@@
At this point, the LTI system’s input parameter matrices are static. Therefore, recent work [16] makes $B$, $C$, and $\Delta$ depend on the input. Recent research suggests that since the $A$ matrix is a HIPPO matrix and $\Delta$ represents the step size, $\exp(\Delta A)$ can be viewed as forget gate and input gate [17], which modulates the influence of the input on the state.
However, traditional SSM lacks the multi-frame querying capabilities that are crucial for BurstSR tasks. Therefore, we propose a QSSM to enable the base frame to gate the output of the current frames, thereby allowing the base frame to perform information queries on the current frames to obtain the sub-pixels while eliminating noise, as illustrated in Figure 1. Specifically, we let the current frames drive the state changes, with the base frame gating the influence of the current frames on the state through $B$ and $\Delta$. As shown in Figure 2 (a), the corresponding formulas are as follows:
@@
Considering that noise may exist at certain locations in the base frame and is randomly distributed at the same locations in other current frames, we concatenate the base frame with other current frames to initially reduce random noise in the original base frame and achieve preliminary feature fusion. Specifically, the concatenated features are processed by an MLP layer and finally projected back to the channel dimension of the original base frame, using it as the new base frame. Then the base frame is transformed through a learnable linear layer to generate $\Delta_{\text{base}_{t}}$ and $B_{\text{base}_{t}}$, and these are then used in the discretization formula from Eq. (2) to obtain $\overline{A}_{\text{base}_{t}}$ and $\overline{B}_{\text{base}_{t}}$. The current frames are processed through another linear layer to obtain $x_{\text{cur}_{t}}$, where $t$ indicates the positional relationship after flattening the base frame and current frames.
$h_{t}&=\sum_{j=0}^{t}\left[\prod_{i=j+1}^{t}\exp(\Delta_{\text{base}_{i}}A)\right]\tilde{f}(\Delta_{\text{base}_{j}})B_{\text{base}_{j}}x_{\text{cur}_{j}},\end{aligned}$
$y_{t}&=\scriptsize{C\sum_{j=0}^{t}\left[\prod_{i=j+1}^{t}\exp(\Delta_{\text{base}_{i}}A)\right]\tilde{f}(\Delta_{\text{base}_{j}})B_{\text{base}_{j}}x_{\text{cur}_{j}}}\\ &\quad+\scriptsize{Dx_{\text{cur}_{t}}},\end{aligned}$
where $\tilde{f}$ represents the functions of $\Delta$ and $B$ corresponding to the zero-order hold of $B$, as follows:
$\tilde{f}(\Delta_{\text{base}_{t}})&=(\Delta_{\text{base}_{t}}A)^{-1}\left(\exp(\Delta_{\text{base}_{t}}A)-I\right)\Delta_{\text{base}_{t}}.\end{aligned}$
[…]
Figure 3: Illustrative comparison between the proposed QSSM and existing cross-attention mechanisms.
Discussion To clarify the differences between the proposed QSSM and existing cross-attention methods, we conduct a detailed analysis along with a schematic comparison. As shown in Figure 3 (a), the cross-attention method uses serialized base feature tokens at position $t$ as the Q matrix, querying with all position tokens of the current features as the K matrix to extract sub-pixel information. In contrast, our QSSM, as shown in Figure 3 (b), not only extracts sub-pixel information from the current features to the base feature but also facilitates information interaction within the base feature. It simultaneously retrieves corresponding current feature tokens from other position tokens in the serialized base feature, thereby achieving joint multi-position sub-pixel information retrieval and noise suppression between the base feature and current features.
3.3 Multi-Scale Fusion Module
Considering the presence of sub-pixel information across various scales within the intricate details of images, we propose a novel Multi-scale Fusion Module (MSFM). This module is designed to effectively integrate multi-scale sub-pixel information from the current frames, thereby enhancing the capability for detailed image reconstruction. The MSFM comprises three distinct branches: a Convolutional Neural Network (CNN), a State Space Model (SSM) with diverse scanning orientations, and a channel Transformer. To begin with, a $3\times 3$ convolution is utilized for the fusion of local sub-pixel features. The SSM is introduced to efficiently learn and integrate sub-pixel features along both horizontal and vertical axes. Furthermore, considering the attenuation characteristics of the A matrix within the SSM when dealing with long-range perception, we concurrently employ a Transformer block to augment the network’s proficiency in capturing global information. The mathematical formulation of the MSFM is as follows:
where $w_{i}$ represents the balancing factors. $x$ and $y$ represent the input feature and output feature, respectively.
3.4 Adaptive Up-sampling Module
After the aforementioned processes, the sub-pixel structural information from burst LR images is extracted and distributed in the feature space. The next critical challenge is to utilize this valuable sub-pixel information to adaptively upsample the image resolution and incorporate sub-pixel details into the high-resolution image. Existing state-of-the-art methods, such as Burstormer [12] and BIPNet [11], simply employ interpolation, transposed convolution, or pixel shuffle techniques for resolution upsampling. However, these methods lack the capability to perceive the distribution of sub-pixels in the feature space, leading to an inability to adaptively reconstruct fine details. Therefore, we introduce a novel Adaptive Up-sampling (AdaUp) module that perceives the distribution of sub-pixels in the spatial domain and adaptively adjusts the up-sampling kernel, thereby achieving higher-quality image detail, as shown in Figure 2 (b). Specifically, we first adaptively perceive the distribution of sub-pixels $L\in\mathbb{R}^{B\times C_{\text{in}}\times 1\times 1}$ from the input features $X\in\mathbb{R}^{B\times C_{\text{in}}\times H\times W}$ by adaptive pooling. We then perform sequence feature interaction on $L$ to obtain the output channel feature distribution sequence $L_{1}\in\mathbb{R}^{B\times C_{\text{out}}\times 1\times 1}$. Subsequently, we apply both the input distribution sequence and the output distribution sequence to the upsampling transposed convolution kernel $W\in\mathbb{R}^{B\times C_{\text{in}}\times C_{\text{out}}\times 3\times 3}$ using broadcasting, thereby endowing the kernel with feature perception capability. Finally, we obtain the high-resolution output through the upsampling transposed convolution kernel. The corresponding formulas are as follows:
$L=\text{AdaptivePooling}(X)$ (9)
$L_{1}=\text{Conv}_{1\times 1}(L)$ (10)
$W_{f}=(W\odot L)\odot L_{1}$ (11)
$y=\text{Trans-Conv}(W_{f},X)$ (12)
where $\odot$ represents element-wise multiplication. Thus, AdaUp can leverage the underlying content information from input frames at different channels and utilize it to get better performance than the mainstream up-sampling operations, pixel shuffle, or interpolations [4, 37].
4 Experiments and Analysis